OpenStudy (anonymous):

What is the solution to g over five equals three over two? 7.5 4 3 10

OpenStudy (anonymous):

7.5

OpenStudy (nurali):

g/5=3/2 Cross multiply 2g=3*5 2g=15 Divide both side by 2 g=?